OpenStudy (anonymous):

Which equation is a point-slope form of the equation of this line? a. -2x+y+5=0 b. y=2x-5 c. (y+1)=-1/2(x-2) d.(y-1)=2(x+2)

OpenStudy (anonymous):



OpenStudy (anonymous):

@Janu16

OpenStudy (princeharryyy):

slope = (7-1)/(1-(-2)) = 2 y1 = 1 x1=-2 required equation => y-y1 = slope*(x-x1) put it there find ur answer

OpenStudy (anonymous):

Im still lost

OpenStudy (princeharryyy):

(x1,y1) = (-2,1) (x2,y2) = (1,7) slope = y2-y1 / x2-x1 point slope form =>y-y1 = slope*(x-x1) u have everything there just insert it in the equation...

OpenStudy (janu16):

ok a and b are off and letter c has the line going other way so its D

OpenStudy (janu16):

@meganbeth1998

OpenStudy (janu16):

wait no its not D

OpenStudy (texaschic101):

its d

OpenStudy (janu16):

wait its is

OpenStudy (texaschic101):

princeharry explained it well

OpenStudy (anonymous):

thanks @Janu16 and @texaschic101 and @princeharryyy

OpenStudy (princeharryyy):

y-1 = 2 (x+2)